Prove that C⁴/C ~_ C³
1
Expert's answer
2020-12-04T06:41:11-0500

Let "\phi:\mathbb{C}^4\rightarrow \mathbb{C}^3" be defined as "\phi(z_1,z_2,z_3,z_4)=(z_1,z_2,z_3)." This projection map is clearly a homomorphism over "\mathbb{C}" , since "\phi(z_1+z_1',z_2+z_2',z_3+z_3',z_4+z_4')=(z_1+z_1',z_2+z_2',z_3+z_3')=(z_1,z_2,z_3)+(z_1'+z_2'+z_3')=\phi(z_1,z_2,z_3,z_4)+\phi(z_1',z_2',z_3',z_4')". Also its clearly, linear since, "\phi(\lambda(z_1,z_2,z_3,z_4))=\phi(\lambda z_1,\lambda z_2,\lambda z_3,\lambda z_4)" "=(\lambda z_1,\lambda z_2,\lambda z_3)=\lambda(z_1,z_2,z_3)=\lambda\phi(z_1,z_2,z_3,z_4)."

Surjectivity is immediate since "(z_1,z_2,z_3)" has "(z_1,z_2,z_3,z_4)" as preimage.

Kernel= "\{(z_1,z_2,z_3,z_4)|\phi(z_1,z_2,z_3,z_4)=(0,0,0,0)\}" Hence Kernel="\{(0,0,0,z_4)\}."

Now let "f:" Kernel"\rightarrow \mathbb{C}" be given by, "f(0,0,0,z_4)=z_4." Clearly, this is an isomorphism. Hence by first isomorphism theorem, "\mathbb{C}^4/Ker \ \phi\cong \mathbb{C}^3\Rightarrow \mathbb{C}^4/Ker \ \phi\cong \mathbb{C}^3."